Solve the system of equations using the elimination method. Enter the answer as an ordered pair.
2x - 3y = 4
5x + 2y = -9

Answer by meetkisai(17)

2(2x-3y) = 4.2
3(5x+2y)=-9.3
4x-6y = 8
15x+6y = -27
19x = -19 => x = -1
2(-1) -3y = 4 = > y = -2
(x, y) => (-1, -2)